Sol Summary – May 18th

Crew 281 Sol Summary Report 18 May 2023
Sol: 4
Summary Title: Rest, research and plan
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
EVA: We did not do any EVA today to give ourselves a break from strenuous activity, focus on the research and get some good planning done for the future EVAs
In the Greenhab, KC (Crew Engineer) and Megan (Commander) established a cross flow technique for passive circulation and investigated potential use of evaporative cooling.
In the science dome, Megan (Commander) has her cacao experiment in progress and Rachel (HSO) worked on her software defined radio.
In the RAM, retrieved supplies for projects in hab and green hab.
Ritupriya (XO) investigated best drone flying techniques to capture scanlines to aid georeferencing and worked with Ana (Crew Scientist) to plan all the future EVA locations and objectives.
After the research and hab work, we did report writing. For dinner we made a tomato sauce based pasta, this sauce was delicious and was made with the green hab harvested tomatoes.
We did some recreational team building and worked on setting up our characters for Dungeons and Dragons.
After 8 pm the crew will be doing a crew bonding exercise.
Look Ahead Plan: Tomorrow we plan on an EVA around the hab to set up ham radio antenna, test rock sampling device, exterior hab inspection and video record the operations
Anomalies in work: None
Weather: Windy, 43 Km/h gusts
Crew Physical Status: No changes
EVA: None
Reports to be filed:
Greenhab Report
Ops Report
Journalist Report
Photo of the Day

Support Requested:
Mission support is requested in case of storms, if the antenna needs to come down or be retrieved.

Journalist Report – May 18th

Crew 281 Journalist Report 18 May 2023

Journalist Report

Crew 281

By: Rachel Jones

Sol 4 was a science day. The team had been pushing hard on EVAs for
the past three days; thus, a break day was needed. The crew slept in
late (i.e., past 0700). Then KC prepared a chocolate chip pancake
brunch from dehydrated scratch. I spent the morning building a
Raspberry Pi system to utilize an RTL-Software Defined Radio (SDR). My
efforts were frustrating as typical troubleshooting methods (i.e.,
looking it up online) were unavailable. This frustration sparked crew
discussion of a future Mars mission’s possible network architecture
and the need for a Mars intranet.

The afternoon brought each individual to pursue their individual
projects. KC, our all-in-one Crew Engineer, performed Hab maintenance
and monitored systems. He also investigated options for optimizing
GreenHab conditions. During the day, that module can get to over 105
*F, which, for some reason, the current varieties of “MDRS” plants
don’t prefer.

Megan checked on her chocolate plants that are currently sprouting in
the Grow Tent in the Science Dome. I got to help her water the
seedlings as I spent my afternoon researching and troubleshooting
SDRs. Thinking of all the parts I wish I had, I realize the ability to
print tools and spare electronic parts on Mars will be necessary, but
for now, I have a few obstacles to overcome.

Ritu and Ana planned the rest of the crew EVAs. Ritu researched drone
flying techniques to capture scan lines. For our mission, a scan line
is the recorded measurements of the geological and geomechanical
characteristics over a specific distance (i.e., the yellow taped line
in yesterday’s photos). Ana also made educational outreach videos for
the Space for All Nations initiative.

Dinner tonight was penne noodles with fresh tomato sauce. Again, this
sauce utilized a mix of vegetables and herbs harvested from the
GreenHab and supplemented with dehydrated vegetables and spices. After
dinner, activities featured writing reports and working on our
Dungeons and Dragons (D&D) characters.

After a day of rest, Crew 281 is ready to go 100% in Sol 5.

GreenHab Report – May 18th

Crew 281 GreenHab Report 18 May 2023
GreenHab Officer: Megan Kane
Environmental control: mixed door open/ door open with fan
Average temperatures: 92.4 °F
Hours of supplemental light: None
Daily water usage for crops: 1.45 gallons
Daily water usage for research and/or other purposes: 0.4 gallons
Water in Blue Tank 172.34 gallons
Time(s) of watering for crops:
0834: refilled 2 water bottle (1 L) watered other plants as needed 1.5 L
1230: refilled 1 water bottle (500 mL) watered other plants as needed 1 L
Changes to crops: None
In the morning the GreenHab was checked and watered as required. Initial temperature was 90.8 degrees Fahrenheit dropping to only 86.9 degrees after the door was propped open. Air was stagnant. As per instructions the fan was not turned on.
Mission control was sent a request to permit working on ideas to cool the GreenHab with the goal of extending the greenhouse operations into the summer weather. This was granted and we began working to improve the airflow. We were able to generate a passive cross breeze in the GreenHab. Supplemental airflow was utilized during the midday peak heat of 105 °F and turned off when the clouds came back.
Utilizing several measurements in the GreenHab and referencing a psychrometric chart, we started the calculations to see if an alternative form of evaporative cooling could be constructed. We did a few tests with materials available to see which would be best for creating an adhoc cooling system. Further work will be done on this tomorrow. The goal being to always keep the GreenHab under 100 degrees. The temperature varied from 86 to 105 degrees Fahrenheit. It is noted that it is hottest in the morning and noon time frame due to the additional shade provided by the hab in the late afternoon.
Harvest: (include which crop and mass in grams): None
Support/supplies needed: None

Operations Report – May 18th

Crew 281 Operations Report 18-05-2023

SOL: 4

Name of person filing report: Megan Kane (Commander)

Non-nominal systems:

Lower deck shopvac filter

Electrical outlets in kitchen

Remote observatory

EVA Helmet

Notes on non-nominal systems:

No new notes.


Rover numbers from previous sol summaries as there was no EVA today.

Spirit rover used: No

Hours: (before EVA): 222.2

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

Opportunity rover used: No

Hours: (before EVA): 129.6

Hours: (after EVA):

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

Curiosity rover used: No

Hours: (before EVA): 239.8

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

Perseverance rover used: No

Hours: (before EVA): 264.3

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

General notes on rovers: no rovers were used

Summary of Hab operations: put notes here

WATER USE: 29.4 gallons

Water (static tank): 396.4 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: yes

Summary of internet:


Summary of suits and radios: All suits and radios are fully recharged.

Summary of GreenHab operations:

WATER USE: 1.85 gallons

Heater: Off

Supplemental light: Off

Harvest: in grams : 0

Summary of Science Dome operations:

Dual split: Heat and AC, Off

Cacao Experiment:

Pulled out and checked cacao seeds

Ham Radio Work:

Worked on software defined radios

Summary of RAM operations: (Tools used, work done)

Retrieved supplies for projects in Hab and Greenhab

Summary of any observatory issues: NA

Summary of health and safety issues:

No new issues to report.

Questions, concerns and requests to Mission Support:

Request Mission Support assistance for the remainder of the mission related to the Ham Radio antenna being installed tomorrow in case of emergency or damage.

Copyright © The Mars Society. All rights reserved. | Main Site